# Copyright in AI Pre-Training Data Filtering: Regulatory Landscape and Mitigation Strategies

### v1

@@ Abstract
The rapid advancement of general-purpose AI models has increased concerns about copyright infringement in training data, yet current regulatory frameworks remain predominantly reactive rather than proactive. This paper examines the regulatory landscape of AI training data governance in major jurisdictions, including the EU, the United States, and the Asia-Pacific region. It also identifies critical gaps in enforcement mechanisms that threaten both creator rights and the sustainability of AI development. Through analysis of major cases we identified critical gaps in pre-training data filtering. Existing solutions such as transparency tools, perceptual hashing, and access control mechanisms address only specific aspects of the problem and cannot prevent initial copyright violations. We identify two fundamental challenges: pre-training license collection and content filtering, which faces the impossibility of comprehensive copyright management at scale, and verification mechanisms, which lack tools to confirm filtering prevented infringement. We propose a multilayered filtering pipeline that combines access control, content verification, machine learning classifiers, and continuous database cross-referencing to shift copyright protection from post-training detection to pre-training prevention. This approach offers a pathway toward protecting creator rights while enabling continued AI innovation.
@@ 1 Introduction
The research community is concerned about the fact that many general-purpose AI (hereinafter GPAI) are trained on data that could not be tracked, meaning that its provenance is unknown. These models are trained on diverse compilations of text, image, and audio data scraped from the web, synthetically generated, or hand-curated [longpre2024data].
@@
This study employs comparative regulatory analysis across the EU, US, and Asia-Pacific jurisdictions, examining copyright frameworks, enforcement cases, and compliance mechanisms such as GDPR [gdpr2016], Copyright Directive [eucopyright2019], AI Act [euaiact2024], DSA [dsa2022] in EU, and case law and state-level legislation in the US. We analyzed lawsuit cases to identify enforcement gaps. We then evaluated existing technical solutions against open problems documented in academic databases such as Stanford’s TAIG Database [stanford2024taig].
Our analysis reveals two critical gaps in current copyright enforcement mechanisms for AI training data that are divided into technical and regulatory. First, pre-training license collection and content filtering emphasizes the impossibility of implementation of copyright clearance mechanisms. Second, the verification mechanisms contain regulatory and technical gaps. These challenges map directly to open questions in Stanford’s Technical AI Governance Database [stanford2024taig]: ”How can licence collection be automated to prevent training on unlicensed data?”, ”How can problematic data be identified without full/direct access to the dataset?”, and ”How could the verification process for correct use of licensed data in AI model training be formalised?”.
@@ 2.1 European Union
In EU AI copyright infringement cases are regulated by EU Copyright Directive [eucopyright2019], while the AI Act [euaiact2024] adds transparency requirements with EU Code of Practice [eu2024codeofpractice] being the guidelines to AI Act. GDPR [gdpr2016] applies to privacy aspects, and DSA [dsa2022] to platform obligations.
2.1.1 EU Code of Practice
In EU Code of Practice, Copyright Chapter, Measure 1.4 [eu2024codeofpractice, Copyright Chapter, Measure 1.4], which is a voluntary framework developed to help AI developers and deployers to comply with the requirements of the EU AI Act, stated that Technical Model Documentation must specify data collection sources, training datasets, and related information to GPAI, however, the details on infringement monitoring systems are not provided. Such systems are critical for ensuring compliance with copyright law and enabling rights holders to identify and address unauthorized use of their work in AI training data. The way that AI Office can monitor infringement cases is through:
1. Technical Model Documentation
AI Office ensures that GPAI providers describe the data filtering methodology used for GPAI in their Technical Model Documentation. This method is based on trust rather than practical surveillance, and the actual audit can be conducted only on case of infringement report.
@@ 2. Reporting Systems
Measure 1.3 of EU AI Code of Practice [eu2024codeofpractice, Copyright Chapter, Measure 1.3] states that the robots.txt protocol should delineate permissible and restricted directories for automated access, and Signatory non-compliance with these access restrictions constitutes a violation of established web crawling conventions and copyright reservation mechanisms. However, if it wasn’t specified that certain directory is not allowed to be accessed, it can be accessed by Signatories. Beyond robots.txt other technical components and restrictions must be respected. This is not limited only to EU content, the restrictions must be respected everywhere. So if OpenAI (US) wanted to train its model on the pirated website (for example operated in Ukraine), it would not be able to sell its services to EU. The EU AI Act has extraterritorial reach, meaning that non-EU companies must comply with all EU AI Code of Practice requirements including copyright restrictions on training data from any global source, if they wish to access the EU market.
@@
Moreover, measure 1.3 states that if an AI Model is also a search engine (for example Google) then the two functions must be separated. For example some website puts ‘no AI training’ in their metadata, but it still wants to be popped up when the user is looking for the information through a search engine. So, Google should not train its Gemini on the data from this website, but still show it to users through the search engine.
@@ 2.1.2 EU Copyright Directive
The Copyright Directive (EU) 2019/790 introduced a general exception allowing AI companies to use copyrighted content for training purposes, but with a critical limitation: rightsholders can ”opt out” by expressly reserving their rights through machine-readable means [eucopyright2019, Article 4(3)]. This creates a legal framework where AI training on copyrighted material is permitted by default, unless the copyright owner has technically implemented an opt-out mechanism on their website or platform. Under Article 4(3), rightsholders must use ”machine-readable means, including metadata and terms and conditions of a website or a service” to reserve their rights against TDM. However, the directive provides limited guidance on what constitutes effective opt-out implementation, creating legal uncertainty for both AI developers and content creators. In Article 3 [eucopyright2019, Article 3] a separate, broader exception exists for research organizations and cultural heritage institutions conducting scientific research, which cannot be overridden by opt-out mechanisms. This creates a two-tiered system where commercial AI development faces restrictions that academic research does not.
@@ 2.1.3 DSA
The DSA mandates creation of a first-of-its-kind regulatory repository where data on content moderation decisions taken by online platforms are accessible to the public at unprecedented scale. This transparency mechanism could potentially track when platforms restrict AI scraping activities or remove content due to unauthorized use in AI training.
The DSA requires very large online platforms (VLOPs) to provide transparency about recommender systems and allow users to opt out of personalized recommendations. The impact of the Digital Services Act on digital platforms — Shaping Europe’s digital future. While not directly addressing AI training, this establishes precedent for algorithmic accountability that could extend to AI training data collection practices.
@@ 2.1.4 GDPR
The GDPR applies to any personal data regardless of its source, requiring controllers to identify a lawful basis under Article 6 before processing personal data for AI training [termly2024gdpr]. Recent enforcement actions, including the Italian DPA’s fine against OpenAI, have focused on companies’ failure to identify appropriate lawful bases for training AI models [dataprotectionreport2025edpb].
Article 14 requires controllers to inform data subjects when processing personal data obtained from sources other than the data subject themselves, including publicly available data [termly2024gdpr]. This creates obligations to respect contextual privacy expectations - for example, scraping should not occur where sites actively prohibit it via robots.txt or from platforms aimed at minors [skadden2025cnil].
Controllers must inform data subjects of their right to object (Article 21) when processing is based on legitimate interests, and in some cases provide specific notification and opportunity to object before processing occurs [dataprotectionreport2024gdpr].
@@ 2.1.5 EU Enforcement Cases
In March 2023, the Italian Data Protection Authority (Garante) launched an investigation into OpenAI’s ChatGPT following concerns about its collection and processing of personal data. The investigation revealed that OpenAI had processed user data to train ChatGPT before the service’s public release on November 30, 2022, without first identifying an adequate legal basis as required under GDPR [LewisSlkin2025OpenAI].
@@
On March 30, 2023, the Garante issued an interim emergency decision ordering OpenAI to immediately stop processing personal data of Italian users through ChatGPT pending further investigation. [CliffordChance2023ChatGPT].
Following discussions between OpenAI and the Garante in April 2023 [Garante2023ChatGPTReinstated], OpenAI proposed remedial measures which were analyzed and approved. On April 28, 2023, the Garante declared that OpenAI had implemented satisfactory measures including a privacy notice for users and non-users, and the right for all individuals in Europe to opt-out from data processing for algorithm training. On December 20, 2024, the Garante published Decision No. 755 (dated November 2, 2024) [DataGuidance2024Garante], imposing a €15 million fine on OpenAI for violations of GDPR Articles 5(1)(a), 5(2), 6, 12, 13, 24, 25, 32, and 33. Beyond the fine, the Garante required OpenAI to conduct a six-month public information campaign via radio, television, print media, and the Internet to inform users and non-users about how ChatGPT works, how data is collected, and how individuals can exercise their rights to object, erasure, and rectification under GDPR.
@@ 2.2 The United States of America
The US lacks the EU’s comprehensive regulatory framework. While multiple lawsuits are pending, there’s no federal legislation providing clear training data restrictions comparable to the EU Copyright Directive’s TDM provisions or the AI Act’s transparency requirements. However, several states within the US are implementing AI-specific copyright transparency requirements. For example, California AB 412 (2025) [california_ab412_2025] requires developers of generative AI models to document copyrighted materials used in training and provide a public mechanism for rights holders to request this information, with 30-day response requirements. Other states are enacting AI legislation, but most focuses on employment discrimination, pricing algorithms, and consumer protection rather than copyright training data restrictions.
@@ 2.2.1 US Enforcement Cases
In December 2023, The New York Times filed a landmark lawsuit against OpenAI and Microsoft in the U.S. District Court for the Southern District of New York, alleging direct copyright infringement [CopyrightAlliance2025Lawsuits, NYTimesComplaint2023]. The complaint alleges that the defendants used millions of the newspaper’s copyrighted articles without permission or compensation to train their large language models, including ChatGPT and Microsoft’s Bing Chat [TechPolicyPress2025ANI]. The training dataset was massive in scale, with models trained on trillions of words—equivalent to a Microsoft Word document over 3.7 billion pages long [HarvardLawReview2024NYT].
The Times argues that this unauthorized use violates copyright law and undermines its business model by enabling users to bypass paywalls and obtain content substantially similar to Times articles without visiting the newspaper’s website [Stempel2023NYTimes]. The lawsuit demonstrates instances where AI models could reproduce content closely resembling specific Times articles, suggesting that training data was being memorized rather than truly transformed [USC2025AICopyright].
OpenAI and Microsoft defend their practices by invoking the fair use doctrine, a uniquely American legal principle that permits limited use of copyrighted material without permission under certain circumstances [USC2025AICopyright]. In their answer to the amended complaint, OpenAI’s attorneys argue that training AI models constitutes ”paradigmatic transformative fair use” because the models learn statistical patterns rather than copying content for redistribution [TheRegister2024OpenAITraining]. They contend that showing copyrighted books to AI models to teach them intelligence and language is analogous to human learning and does not constitute infringement because it extracts only word frequencies, syntactic patterns, and other statistical data rather than reproducing the works themselves [TheRegister2024OpenAITraining].
In November 2024, the court issued a significant ruling denying OpenAI’s motion to compel evidence related to the New York Times’ business practices or its employees’ use of generative AI [CopyrightAlliance2025Lawsuits]. The court determined that such discovery was irrelevant to the fair use analysis, rejecting OpenAI’s attempt to argue that the Times’ own use of AI undermined its claims. This ruling clarified that a fair use defense must stand on its own merits rather than on the plaintiff’s behavior.
The case has been consolidated with other lawsuits filed by news organizations, including Daily News et al. v. Microsoft Corp. et al. and The Center for Investigative Reporting v. OpenAI, creating a unified proceeding that will address copyright issues affecting the entire news publishing industry [CopyrightAlliance2025Lawsuits].
@@ 2.3 Comparison of EU and US approaches
Unlike the EU’s proactive regulatory enforcement through GDPR and the AI Act, the United States lacks comprehensive federal legislation governing AI training data [USC2025AICopyright]. The US approach is characterized by:
@@
    Case-by-case adjudication: Rather than establishing clear rules, the US legal system relies on courts to determine whether AI training constitutes fair use through individual lawsuits [CopyrightAlliance2025Lawsuits].
@@
    Industry self-regulation: In the absence of federal mandates, AI companies have not been required to disclose training datasets or implement filtering systems, though some states are beginning to act [USC2025AICopyright].
@@
    Emphasis on fair use balancing: US courts will weigh four statutory factors: (1) purpose and character of use, (2) nature of the copyrighted work, (3) amount used, and (4) effect on the market for the original work [USC2025AICopyright].
The New York Times case exemplifies the reactive, litigation-based US approach to AI governance, which contrasts sharply with the EU’s proactive regulatory model. While the EU imposes immediate compliance obligations through GDPR and the AI Act—including transparency requirements, legal basis documentation, and substantial fines for violations—the US system allows AI companies to operate without such constraints while legal battles slowly proceed through the courts. The outcome of this case could fundamentally reshape how AI models are trained in the United States, potentially requiring licensing agreements with content creators or forcing companies to rely solely on public domain or licensed materials [USC2025AICopyright].

### v2

@@ Abstract
The rapid advancement of general-purpose AI models has increased concerns about copyright infringement in training data, yet current regulatory frameworks remain predominantly reactive rather than proactive. This paper examines the regulatory landscape of AI training data governance in major jurisdictions, including the European Union (or EU), the United States (or US), and the Asia-Pacific region. It also identifies critical gaps in enforcement mechanisms that threaten both creator rights and the sustainability of AI development. Through analysis of major cases we identified critical gaps in pre-training data filtering. Existing solutions such as transparency tools, perceptual hashing, and access control mechanisms address only specific aspects of the problem and cannot prevent initial copyright violations. We identify two fundamental challenges: pre-training license collection and content filtering, which faces the impossibility of comprehensive copyright management at scale, and verification mechanisms, which lack tools to confirm filtering prevented infringement. We propose a multilayered filtering pipeline that combines access control, content verification, machine learning classifiers, and continuous database cross-referencing to shift copyright protection from post-training detection to pre-training prevention. This approach offers a pathway toward protecting creator rights while enabling continued AI innovation.
@@ 1 Introduction
The research community is concerned about the fact that many general-purpose AI (hereinafter GPAI) models are trained on data that could not be tracked, meaning that its provenance is unknown. These models are allegedly trained on diverse compilations of text, image, and audio data scraped from the web, synthetically generated, or hand-curated [16].
@@
This study employs comparative regulatory analysis across the EU, US, and Asia-Pacific jurisdictions, examining copyright frameworks, enforcement cases, and compliance mechanisms such as the General Data Protection Regulation (hereinafter GDPR) [gdpr2016], Copyright Directive [eucopyright2019], AI Act [euaiact2024], the Digital Services Act (hereinafter DSA) [dsa2022] in EU, and case law and state-level legislation in the US. We analyzed lawsuit cases to identify enforcement gaps. We then evaluated existing technical solutions against open problems documented in academic databases such as Stanford’s Technical AI Governance (hereinafter TAIG) Database [22].
Our analysis reveals two critical gaps in current copyright enforcement mechanisms for AI training data that are divided into technical and regulatory. First, pre-training license collection and content filtering emphasizes the impossibility of implementation of copyright clearance mechanisms. Second, the verification mechanisms contain regulatory and technical gaps. These challenges map directly to open questions in Stanford’s Technical AI Governance Database [22]: ”How can licence collection be automated to prevent training on unlicensed data?”, ”How can problematic data be identified without full/direct access to the dataset?”, and ”How could the verification process for correct use of licensed data in AI model training be formalised?”.
@@ 2.1 European Union
In the EU, copyright infringement cases involving AI-generated content fall under the EU Copyright Directive [eucopyright2019]. The EU AI Act [euaiact2024] introduces transparency obligations for providers of general-purpose AI (GPAI) models. The GDPR [gdpr2016] applies to privacy aspects, while the DSA [dsa2022] to platform obligations.
2.1.1 EU Code of Practice for GPAI Model Providers
The EU Code of Practice for GPAI model providers serves as a voluntary framework designed to help GPAI model providers comply with the requirements of the EU AI Act. The Copyright Chapter, Measure 1.4 [eu2024codeofpractice, Copyright Chapter, Measure 1.4] suggests to mitigate the risk of copyright-infringing outputs. The way that AI Office can monitor infringement cases is through:
1. Model Documentation
GPAI model providers are required to describe the data filtering methodology used for GPAI in their Model Documentation. This method relies exclusively on trust rather than technical verification, and such verifications are carried out only when an infringement is reported or detected.
@@ 2. Reporting Systems
Different information may be required from the GPAI model providers and it differs depending on the recipient, for example, AI Office, downstream provider, national competent authority.
Measure 1.3 of EU AI Code of Practice [eu2024codeofpractice, Copyright Chapter, Measure 1.3] states that the robots.txt protocol should delineate permissible and restricted directories for automated access, and Signatory non-compliance with these access restrictions constitutes a violation of established web crawling conventions and copyright reservation mechanisms. Beyond robots.txt other technical components and restrictions must be respected. This is not limited only to EU content, the restrictions must be respected everywhere. So if OpenAI (US) wanted to train its model on a pirated website (for example operated in Ukraine), it would not be able to sell its services to the EU content, as the providers should not train their model on any illegal content (Measure 1.2 of Copyright Chapter “Reproduce and extract only lawfully accessible copyright-protected content when crawling the World Wide Web”). The EU AI Act has extraterritorial reach, meaning that non-EU companies must comply with all EU AI Code of Practice requirements including copyright restrictions on training data from any global source, if they wish to access the EU market.
@@
Moreover, measure 1.3 states that if a GPAI model provider also provides a search engine (for example Google) then the two functions must be separated. For example, some website puts ‘no AI training’ in their metadata, but it still wants to be referenced when the user is looking for the information through a search engine. So, Google should not train its Gemini on the data from this website, but still show it to users through the search engine.
@@ 2.1.2 EU Copyright Directive
The Copyright Directive (EU) 2019/790 introduced exceptions for Text and Data Mining (TDM) which is the automated analysis of digital content to extract patterns, trends, and insights. Article 4 provides a general TDM exception that some interpret as potentially allowing AI companies to use copyrighted content for training purposes, though whether this exception applies to AI model training remains legally uncertain [8]. This provision includes a critical limitation: rightsholders can ”opt out” by expressly reserving their rights through machine-readable means [eucopyright2019, Article 4(3)]. This creates a legal framework where TDM on copyrighted material is permitted by default, unless the copyright owner has technically implemented an opt-out mechanism on their website or platform. Under Article 4(3), rightsholders must use ”machine-readable means, including metadata and terms and conditions of a website or a service” to reserve their rights against TDM. However, the directive provides limited guidance on what constitutes effective opt-out implementation, creating legal uncertainty for both AI developers and content creators. In Article 3 [eucopyright2019, Article 3] a separate, broader exception exists for research organizations and cultural heritage institutions conducting scientific research, which cannot be overridden by opt-out mechanisms. This creates a two-tiered system where commercial AI development faces potential restrictions that academic research does not.
@@ 2.1.3 DSA
The DSA mandates the creation of a first-of-its-kind public repository of content moderation decisions made by online platforms. This transparency mechanism could help track when platforms restrict AI scraping activities or remove content due to unauthorized use in AI training.
The DSA requires very large online platforms (VLOPs) to provide transparency about recommender systems and allow users to opt out of personalized recommendations. While not directly addressing AI training, this establishes precedent for algorithmic accountability that could extend to AI training data collection practices.
@@ 2.1.4 GDPR
The GDPR applies to the processing of personal data within its territorial scope, regardless of whether the data is obtained directly from data subjects or from publicly available or third-party sources. [27]. Recent enforcement actions, including proceedings involving OpenAI, have focused on inadequate justification of lawful bases, particularly failures in interest balancing, transparency, and purpose limitation in the context of AI training.[6].
Article 14 requires controllers to inform data subjects when processing personal data obtained from sources other than the data subject themselves, including publicly available data [27]. This creates obligations to respect contextual privacy expectations - for example, scraping should not occur where sites actively prohibit it via robots.txt or from platforms aimed at minors [20].
Controllers must inform data subjects of their right to object (Article 21) when processing is based on legitimate interests, and in some cases provide specific notification and opportunity to object before processing occurs [5].
@@ 2.1.5 EU Enforcement Cases
In March 2023, the Italian Data Protection Authority (Garante) launched an investigation into OpenAI’s ChatGPT following concerns about its collection and processing of personal data. The investigation revealed that OpenAI had processed user data to train ChatGPT before the service’s public release on November 30, 2022, without first identifying an adequate legal basis as required under GDPR [13].
@@
On March 30, 2023, the Garante issued an interim emergency decision ordering OpenAI to immediately stop processing personal data of Italian users through ChatGPT pending further investigation. [2].
Following discussions between OpenAI and the Garante in April 2023 [9], OpenAI proposed remedial measures which were analyzed and approved. On April 28, 2023, the Garante declared that OpenAI had implemented satisfactory measures including a privacy notice for users and non-users, and the right for all individuals in Europe to opt-out from data processing for algorithm training. On December 20, 2024, the Garante published Decision No. 755 (dated November 2, 2024) [7], imposing a €15 million fine on OpenAI for violations of GDPR Articles 5(1)(a), 5(2), 6, 12, 13, 24, 25, 32, and 33. Beyond the fine, the Garante required OpenAI to conduct a six-month public information campaign via radio, television, print media, and the Internet to inform users and non-users about how ChatGPT works, how data is collected, and how individuals can exercise their rights to object, erasure, and rectification under GDPR.
@@ 2.2 The United States of America
The US lacks the EU’s comprehensive regulatory framework. While the EU Copyright Directive includes text and data mining (TDM) exceptions, their applicability to large-scale AI model training remains legally contested. In contrast, U.S. defendants rely on the open-ended fair use doctrine, which, like TDM, may be invoked to justify training activities, but is not a purpose-built exception for AI. However, several states within the US are implementing AI-specific copyright transparency requirements. Several U.S. states have proposed AI-specific copyright transparency obligations. For example, California Assembly Bill 412 (2025) [1] would have required developers of generative AI systems to document copyrighted materials used in training and to provide disclosure mechanisms for rights holders. However, the bill was not adopted, reflecting ongoing political and legal resistance to imposing training-data transparency obligations at the state level. Other states are enacting AI legislation, but most focus on employment discrimination, pricing algorithms, and consumer protection rather than copyright training data restrictions.
@@ 2.2.1 US Enforcement Cases
In December 2023, The New York Times filed a landmark lawsuit against OpenAI and Microsoft in the U.S. District Court for the Southern District of New York, alleging direct copyright infringement [4, 29]. The complaint alleges that the defendants used millions of the newspaper’s copyrighted articles without permission or compensation to train their large language models, including ChatGPT and Microsoft’s Bing Chat [26]. The training dataset was massive in scale, with models trained on trillions of words—equivalent to a Microsoft Word document over 3.7 billion pages long [10].
The Times argues that this unauthorized use violates copyright law and undermines its business model by enabling users to bypass paywalls and obtain content substantially similar to Times articles without visiting the newspaper’s website [23]. The lawsuit demonstrates instances where AI models could reproduce content closely resembling specific Times articles, suggesting that training data was being memorized rather than truly transformed [12].
OpenAI and Microsoft defend their practices by invoking the fair use doctrine, a uniquely American legal principle that permits limited use of copyrighted material without permission under certain circumstances [12]. In their answer to the amended complaint, OpenAI’s attorneys argue that training AI models constitutes ”paradigmatic transformative fair use” because the models learn statistical patterns rather than copying content for redistribution [30]. They contend that showing copyrighted books to AI models to teach them intelligence and language is analogous to human learning and does not constitute infringement because it extracts only word frequencies, syntactic patterns, and other statistical data rather than reproducing the works themselves [30].
In November 2024, the court issued a significant ruling denying OpenAI’s motion to compel evidence related to the New York Times’ business practices or its employees’ use of generative AI [4]. The court determined that such discovery was irrelevant to the fair use analysis, rejecting OpenAI’s attempt to argue that the Times’ own use of AI undermined its claims. This ruling clarified that a fair use defense must stand on its own merits rather than on the plaintiff’s behavior.
The case has been consolidated with other lawsuits filed by news organizations, including Daily News et al. v. Microsoft Corp. et al. and The Center for Investigative Reporting v. OpenAI, creating a unified proceeding that will address copyright issues affecting the entire news publishing industry [4].
@@ 2.3 Comparison of EU and US approaches
Unlike the EU’s proactive regulatory enforcement through GDPR and the AI Act, the United States lacks comprehensive federal legislation governing AI training data [12]. The US approach is characterized by:
@@
    Case-by-case adjudication: Rather than establishing clear rules, the US legal system relies on courts to determine whether AI training constitutes fair use through individual lawsuits [4].
@@
    Industry self-regulation: In the absence of federal mandates, AI companies have not been required to disclose training datasets or implement filtering systems, though some states are beginning to act [12].
@@
    Emphasis on fair use balancing: US courts will weigh four statutory factors: (1) purpose and character of use, (2) nature of the copyrighted work, (3) amount used, and (4) effect on the market for the original work [12].
The New York Times case exemplifies the reactive, litigation-based US approach to AI governance, which contrasts sharply with the EU’s proactive regulatory model. While the EU imposes immediate compliance obligations through GDPR and the AI Act—including transparency requirements, legal basis documentation, and substantial fines for violations—the US system allows AI companies to operate without such constraints while legal battles slowly proceed through the courts. The outcome of this case could fundamentally reshape how AI models are trained in the United States, potentially requiring licensing agreements with content creators or forcing companies to rely solely on public domain or licensed materials [12].